geerky42 (geerky42):

You have a 100 pound bag of potatoes. 99% of the bag is water weight. You leave the bag in the sun for a few hours. Now you bring it in and it is only 98% water weight, due to evaporation. What is the new total weight of the bag of potatoes?

geerky42 (geerky42):

@amistre64 @Hero @jayz657 @cwrw238

OpenStudy (anonymous):

im not exactly sure but it might be just 98 pounds (0.98)(100) I could be wrong, need someone to double check

OpenStudy (anonymous):

Yeah i was going to say the same, but it seems too simple for that to be the answer

geerky42 (geerky42):

I'm not sure. If bag is made up to 99% of water, then it must contains one pound of dry potatoes. Now, it contains 98% of water weight, I am not sure how to handle this nor what to do next...

geerky42 (geerky42):

@surdawi

OpenStudy (cwrw238):

you could do this using algebra suppose amount of water lost = x pounds weight of water after being in sun = 99 - x weight of potatoes = 100 - x 98% of this is water so 99-x ----- = 0.98 100 - x find x 100-x will be the new weight of potatoes

geerky42 (geerky42):

It just sound so wrong, but I got 50... I don't think my answer is correct.

OpenStudy (cwrw238):

hmmm - see what you mean - the potatoes have lost half their weight!!! doesn't seem to make sense....

geerky42 (geerky42):

Well, it seems that we didn't do math correctly. I don't know, lol...

OpenStudy (cwrw238):

50 is the correct solution of the equation - you were correct maybe the equation is wrong....

OpenStudy (cwrw238):

the potatoes would have shriveled up to half their weight

OpenStudy (cwrw238):

but i can't see anything wrong with my logic i know i'm tired !!!!! lol

OpenStudy (cwrw238):

50 pounds have evaporated in a few hours in the sun - i suppose its possible if they are in the arizona or saudi dessert

OpenStudy (cwrw238):

i think its the correct answer @geerky42

OpenStudy (cwrw238):

@juanjohnguy - can you see where i might be wrong . i've made human errors before as we all have.

OpenStudy (anonymous):

I was actually just looking at your logic so let me think it through a little

OpenStudy (anonymous):

The math seems right although i seem to be struggling with why you think that you can make those equations for it

OpenStudy (cwrw238):

ok you lose x pounds of water in the sun that means there is 99-x pounds of water left in the potatoes and the new weight of the potatoes is 100-x pounds 98% or 0.98 of the 'new' potatoes i now water so ( 99-x )/ (100-x) = 0.98

OpenStudy (cwrw238):

i can't see anything wrong with that reasoning - but its a surprising result but i'm getting more convinced that its correct

OpenStudy (anonymous):

Yeah it actually seems correct and I'm not finding a way to disprove it

OpenStudy (cwrw238):

its a bit like the question about the rope tied tightly around the earth - if you lengthened it by 3 feet how high would it be above the ground?

OpenStudy (cwrw238):

- the answer is 6 inches